What is the energy in joules and eV of a photon in
Answer #1

  Energy of a photon = h f,
where h = Planck's constant = 6.63 × 10-34 J / s
and f = frequency = 1385 x 10^6 Hz.
=> energy of a photon
= (6.63 × 10^-34) * (1385 x 10^6) J
= 9.18 x 10^-25 J.

Energy of photon in 1385 Mhz FM radio wave
= hf = 6.63x10^-34x1385x10^6/(1.6x10^-19) = 5.74x10^-6 eV